If the [H+] of a solution is 5.7 x 10-14 , what is the [OH-]? Round to four decimal places.


1
Expert's answer
2021-04-19T05:07:42-0400

"Kw=[H+][OH-]"

"Where Kw=1\times 10^{-14}"

Therefore:

"OH- =\frac{Kw}{H+}"

"OH- = \frac{1\times 10^{-14}}{5.7\times10^{-14}}=0.1754"